# Rise and Fall of Anderson Localization By Lattice Vibrations: A Time-Dependent Machine Learning Approach

### v1

@@ III Results
As a real-world example of electron-lattice dynamics, at least within a Fröhlich Hamiltonian, we investigate the prototypical strange metal lanthanum strontium copper oxide (LSCO), renowned for its diverse physics [43]. This material has a large electron-lattice coupling, and its Fermi energy is adjustable through doping. The material parameters for optimal doping are given in Table 1 based on experimental values derived from Refs. [44, 45, 46, 47, 48]. These serve as the basis for constructing an associated deformation potential. These parameters align with previous investigations on strange metals  [14, 13] using a quantum-acoustical perspective. Here we explore electron-lattice dynamics in a broader scope, rather than focusing on specific attributes like electrical conductivity.
To enable this analysis, we introduce two scaled variables that we vary in our simulations: dimensionless temperature $\tilde{T}=T/T_{D}$ and effective electron-lattice coupling $\tilde{G}=2k_{F}/q_{D}$. The coupling is adjusted by varying the Fermi wavevector $k_{F}$ (energy) of the electron while maintaining the underlying lattice structure constant (with a fixed Debye wavevector $q_{D}$). This ensures that our variables $\tilde{T}$ and $\tilde{G}$ are independent, a premise supported by the evidence from our simulations.
@@ III.1 Phase Diagram
Our central finding is presented in Fig. 3 showing the dynamical data classified using the ML-based clustering algorithm explained above, and as detailed in Appendix A as the temperature $\tilde{T}$ and coupling strength $\tilde{G}$ are varied. Three distinct phases are identified, as labeled by the differently colored regions. We want to emphasize that the term “phase” is used here to refer to regimes of different dynamical behavior, not in the thermodynamical sense. There are no sharp boundaries between these phases; the changes are gradual, rather than true phase transitions. This fact is highlighted in Fig. 3 by the different sizes of the points, which represent the level of agreement within the ensemble of studied wavepackets for the given parameters.
Figure 3: Phase diagram of LSCO in the dynamic potential field. The phase diagram was derived using a machine learning-based clustering algorithm to analyze time series data of the wavefunction evolution within systematically varied deformation potentials. This analysis involved variations in temperature $\tilde{T}$ and effective coupling $\tilde{G}$. The following three clusters were identified: I) refractive scattering region (green), II) diffraction behavior (blue), and III) short-time localization at high temperatures (red). The size of the points indicates the level of agreement across an ensemble of different wavefunction data sets.
@@
Phase I (green) is characterized by an almost linear phase boundary starting at $\tilde{T}\approx 0.45$ rising across the range of $\tilde{G}$ explored. This phase is perturbative in the sense of $\tilde{K}\gg 1$. As seen in the left column of Fig. 4, the scattering of the wavepacket is mainly refractive. This trend will lead to branched flow behavior [34] at longer times; a propagating wave forms tree-like branches under a weakly disordered medium, due to small-angle refraction [49]. Moreover, there is a partial transparency of the electrons to any shorter wavelength modes ($q>2k_{F}$) present in the underlying deformation potential, as is further explained in Ref. [10].
Phase II (blue) covers the upper right section of the phase diagram with high values of temperature variability ($\tilde{T}$) and effective coupling($\tilde{G}$) and is exponentially separated by a phase boundary from Phase III (red), which is characterized by high temperatures but lower $\tilde{G}$ levels. Like Regime I, the second phase is characterized by relatively adiabatic lattice dynamics. In other words, the deformation landscape appears as if it is stationary for an electron, at least for short times of $\sim 2\pi/\omega_{D}$. This fact is further confirmed by our IPR results below. Furthermore, this regime is perturbative but also classical-like, meaning that the wavelength of the electron is shorter than the effective shortest length scale of the deformation potential. As thoroughly discussed in Ref. [10], the perturbation theory pathway, particularly Fermi’s golden rule, is proven to be highly successful in this phase.
@@
The final phase (Phase III) identified by the ML-clustering is associated with highly nonperturbative ($\tilde{K}\lesssim 1$) electron-lattice interaction, primarily existing in the parameter space where electron dynamics can be considered as coherent ($\tilde{\lambda}\lesssim 1$). Therefore, wave interference and diffraction effects are important because the electron wavelength is larger than the shortest length scale of the deformation potential. Notably, this phase begins at low temperature as $\tilde{T}\approx 0.5$, while extending to very high temperatures, across the range of $\tilde{T}$ investigated.
In Phase III, as depicted in the right column of Fig. 4, an initial wavepacket encounters significant scattering from a strong deformation potential, initially causing diffusive behavior akin to that seen in Phase II. However, wavepacket spreading eventually ceases due to quantum interference effects, signifying an onset of localization. Nevertheless, the random fluctuations introduced by the motion of the lattice slowly but surely scramble the quantum interference required for the long-term confinement of the wavepacket, resulting in the transient nature of this localization (for capturing the essential aspects of this phenomenon, see, e.g., Ref. [9]).
To achieve a more comprehensive understanding, we adopt a static potential approximation, wherein the temporal aspect of the lattice deformation field of Eq. 2 is neglected, effectively frozen into its original configuration. Within this frozen potential framework, we carry out an analysis analogous to that of the evolving deformation potential as above. This is outlined in Appendix B. Employing the same cluster classification, we categorize the data of an electron wavepacket evolution under a static deformation potential, yielding a phase diagram similar to its dynamic counterpart in Fig. 3, albeit with slightly sharper phase boundaries. This comparison validates treating the deformation potential as predominantly static, particularly in Phases I and II. On the contrary, in Phase III, the static deformation potential results in full Anderson localization of the wavepacket that proves transient when the deformation potential undergoes morphing and undulation over time, as further elucidated in the subsequent analysis.
@@ III.2 Transient Localization
In this section, we delve deeper into the nature of transient localization induced by lattice vibrations taking place within Phase III. At timescales shorter than the characteristic timescale of $2\pi/\omega_{D}\sim 100\,\textrm{fs}$, lattice vibrations mimic a static, internal disorder field, triggering the onset of Anderson localization. Therefore, in reference to the dynamical field, where the motion of the lattice disrupts the process of Anderson localization, we also explore the localization behavior of a wavepacket within a frozen potential approximation. In both cases, we quantify the level of localization by studying the time-dependent IPR, denoted as $\alpha(t)$ in Eq. (7), and introducing a subsidiary product of the MSD, called the instantaneous diffusivity
@@
where $\beta(t)$ is defined in Eq. (6).
Figure 5: Inverse scattering rate $1/\tau$ for electrons in LSCO in a (a) frozen and (b) dynamic deformation potential on a normalized scale. Both scenarios exhibit a high inverse scattering rate region in the upper left corner (Phase III), suggesting strong constraints on carrier mobility. This region is markedly more prominent in the frozen potential scenario, as the lattice vibrations in the dynamic potential continuously disrupt short-lived attempts at localization. The shape of counter lines closely resembles the dynamical phase transition lines depicted in Fig. 4.
@@
consistent with the definition used in Ref. [13]. Figure 5 shows the inverse scattering rate $1/\tau$ for both the cases of frozen (left panel) and dynamic (right panel) deformation potential. Overall, the analysis of the scattering rate supports the ML classification underlying the phase diagram shown in Fig. 4. Both scenarios exhibit a notably high inverse scattering rate within Phase III of the phase diagram (upper right corner), indicating significant constraints on carrier mobility, as expected in the context of Anderson localization. This effect is more prominent in the frozen potential approximation than in the case of the morphing deformation potential, underlining the fact that the dynamics of the deformation potential continuously disrupt short-lived localization attempts. Moreover, we observe that the contour lines of the inverse scattering time in Fig. 5 closely resemble the phase boundaries seen in Fig. 4.
@@
A deeper insight into the emergence of transient localization is obtained by investigating the IPR of the wavepacket evolving over time. In the left panel of Fig. 6, we show the evolution of the IPR in Phase III for both static (blue) and dynamic (green) deformation potentials. In the approximation where the deformation potential remains static at its initial state, the IPR stabilizes at a certain value ($\beta\sim 0.3$) after an initial decrease, heralding the Anderson localization of the wavepacket. Similarly, when the wavepacket is subject to a dynamic deformation potential, a form of Anderson localization occurs, with slightly stronger localization observed compared to the frozen potential case, identified to be caused by small potential variations. This localized state is eventually disrupted by lattice motion, leading to a temporary delocalization followed by a brief re-localization before being disintegrated again by potential evolution. This cyclical process of plateauing IPR seen in the left panel of Fig. 6 epitomizes the birth and demise of the Anderson localization due to lattice vibrations.
Figure 6: Normalized inverse participation ratio $\beta(t)$ of the wavepacket is analyzed as a function of time across three distinct dynamical regions identified in the machine learning-based phase diagram shown in Fig. 4. Figure contrasts the behaviors under static (blue line) and dynamic (green line) deformation potentials. In Phase III (left), brief localization periods (indicated by dashed lines) are observed due to lattice motion, while full localization, indicated by the saturation of $\beta(t)$, occurs within the frozen potential approximation. In Phase II (middle) and Phase I (right), the decay of $\beta(t)$ follows a relatively exponential trend without signs of localization as observed in Phase III. Notably, the decay is faster in Phase I compared to Phase II, where small oscillations linked to quantum interference are present
@@
We would like to point out that the dynamics of short wavelength components of the deformation potential can significantly influence localization within a time window shorter than the characteristic time $\omega_{D}/2\Pi$. For example, the left panel of Fig. 6 demonstrates an instance of dynamically enhanced localization: The random initial configuration of the potential creates valleys (mountains) that quickly move towards (away) the wavefunction, causing boosted localization. Vice versa, these types of small potential variations in time can also lead to weaker localization compared to the static potential case in the first localization plateau seen in the left panel of Fig. 6.
@@
To provide the full picture, we also present the evolution of the IPR measure for Phase I and II in the right and left panels of Figure 6, respectively. Neither phase exhibits any signs of localization, contrasting with the behavior observed in Phase III. Moreover, the overlapping curves of the static and dynamic potentials further support the earlier assertion that the deformation potential can be effectively approximated as a static entity in Phase I and II. In Phase II, the IPR shows rapid exponential decay, quickly approaching the ergodic (fully delocalized) limit of $\beta\sim 0$. This behavior resembles that of a system with a high density of impurities characterized by Gaussian statistics. Similarly, in Phase I, the IPR decreases towards the ergodic limit, albeit at a slower rate, displaying subtle oscillation. The slower, non-exponential decay can be attributed to the weak, refractive nature of wavepacket scattering, in conjunction with quantum coherence and interference effects.
@@ IV Conclusion and Future Directions
Quantum acoustics open up an unexplored pathway to investigating the intricate Fröhlichian electron-lattice interaction, inaccessible with the standard methods of perturbation theory. We take the concept further by treating it not just as a dual perspective on lattice vibrations, but as a versatile tool in ascendance: a time-dependent, nonperturbative approach for electron-lattice interaction in coordinate space. Moreover, in the quasi-classical limit of the coherent state formalism, the quantum-acoustical way unveils the dynamics of electrons navigating through an internal lattice disorder field undulating and propagating in time.
@@
Our study, supported by machine learning, explores the parameter space characterized by temperature and effective coupling, focusing on the paradigmatic strange metal LSCO, known for its two-dimensional transport behavior. However, the presented method can be readily extended to variations in any set of material parameters – potentially augmented by density functional simulations as they are not necessarily independent – and generalized to electron-lattice dynamics in three dimensions. Therefore, our work not only lays the groundwork for uncovering hidden realms in the electron-lattice interaction but is also a testament for designing materials with customized features by employing machine learning techniques from a point of view of dynamics.
@@ Appendix A Clustering Analysis
    Update Step: Update the centroids using $\mathcal{S}$ according to Eq. (10).
It is easy to see that these steps cannot increase the in-set variance – the algorithm is guaranteed to converge. However, there is an important caveat in that the objective function $\mathcal{L}(\mathcal{S})$ defined in Eq. (9) is non-convex and the algorithm can therefore converge to a local optimum that is not the global minimum. This problem can be mitigated by an intelligent choice of initialization scheme, most commonly $k$-means++ [51], and by running the algorithm multiple times and taking the clustering with minimal $\mathcal{L}(\mathcal{S})$ post convergence [16]. An example run of $k$-means on toy data in $\mathbb{R}^{2}$ is shown in Fig. A.1.
Figure A.1: Toy example to illustrate a run of the $k$-means algorithm with hyperparameter $k=3$. (a) Initial centroids are chosen by an initialization scheme (here: random initialization). (b) Assignment step of the first iteration. (c) Update step of the first iteration. (d) Final result. The algorithm converges after two consequent iterations yield sufficiently similar clusters (here: $10$th iteration).
@@
When working with time-series data $x(t)=(x^{t_{1}},x^{t_{2}},\ldots,x^{t_{M}})$, $x^{t_{i}}\in\mathbb{R}^{d}$, one could naively embed $x(t)$ in the space $\mathbb{R}^{M\times d}$ equipped with the Euclidean metric and run $k$-means exactly as described above. However, this approach has significant limitations: It only works for sequences of equal length, and more critically, the Euclidean metric is unable to account for the temporal nature of the sequences. By operating solely on elements with the same time indices, it ignores potential temporal misalignment between sequences. Consequently, even time series with similar features can result in a large metric distance if their phases differ. This makes it an ineffective measure for determining similarity. Dynamic time warping (DTW), introduced by Sakoe and Chiba [42], has been employed to overcome these challenges by identifying the most temporally appropriate pairs of time indices for comparison [52]. Formally,
@@
where the time index pairs $\pi=\{\pi_{0},\pi_{1},\ldots,\pi_{K}\}$ satisfy certain (boundary) conditions that we will not elaborate on for brevity (for an extensive introduction to the topic, see, e.g., Ref. [53]). One thus ends up with a new warped time path for the comparison of the time series, as is illustrated in Fig. A.2.
Figure A.2: Illustration of the difference between the Euclidean distance and Dynamic Time Warping (DTW) distance for two example MSD sequences from our data set. While the Euclidean distance compares values at corresponding time steps, DTW temporally aligns the sequences, providing a more robust distance measure for time-series data.
DTW is not a metric, because $\text{DTW}(x(t),y(t))=0$ does not imply $x(t)=y(t)$ and it doesn’t obey the triangle inequality. Nevertheless, the Fréchet mean, which generalizes Eq. (10) to any metric space, is used to compute a centroid-like representation, commonly referred to as barycenter:
@@
$\sigma_{t}=\sqrt{\frac{1}{T}\int_{0}^{T}\left(x(t)-\mu_{t}\right)^{2}\,\text{d}t}.$
@@
Figure A.3: Elbow plot depicting the sum of squared distances of samples to their nearest cluster center for varying values of $k$ for the (a) frozen and (b) dynamic lattice vibration field scenario. The elbow point, where the curve shows a distinct bend, suggests the optimal number of clusters for our analysis, as increasing the number of clusters beyond this value yields diminishing returns in terms of model improvement.
The parameter $k$, i.e. the number of clusters, has to be set before performing the optimization Eq. (9). To find a suitable $k$ for our analysis, we run the $k$-means algorithm multiple times with varying values of $k$ ranging from 1 to 10. Fig. A.3 shows a plot of the sum of squared distances of samples to their nearest cluster center (inertia) against different values of $k$. The point where the curve bends or forms an elbow is considered an indicator of the natural number of clusters in the underlying data [16] – additional clusters beyond this point do not substantially improve the model’s representation of the different dynamical regimes of the system.
@@
To mitigate the effect of statistical fluctuations of the lattice vibration field on the clustering of the electron dynamics, we perform the optimization independently on an ensemble of 10 wavefunction data sets, each generated using deformation potentials with different random initializations. The different clustering results are then consolidated into one phase diagram by taking the consensus of the ensemble, with the size of the dots in Fig. 3 indicating the degree of consensus.

### v2

@@ III Results
As a real-world example of electron–lattice dynamics, at least within the Fröhlich Hamiltonian, we investigate the prototypical strange metal lanthanum strontium copper oxide (LSCO), renowned for its diverse physics [52]. This canonical cuprate, discovered by Bednorz and Müller [53], is characterized by an orthorhombic space group [54]. The main electrical transport occurs between the $\textrm{CuO}_{2}$ layers, making it effectively two-dimensional in nature [52]. Furthermore, LSCO has a large electron–lattice coupling, and its Fermi energy is adjustable through doping. The material parameters for optimal doping are given in Table 1 based on experimental values derived from Refs. [55, 56, 57, 58, 59]. These serve as the basis for constructing an associated deformation potential. These parameters align with previous investigations on strange metals [25, 24] using a quantum-acoustical perspective. Here, we explore electron–lattice dynamics in a broader scope, rather than focusing on specific attributes like electrical conductivity.
To enable this analysis, we introduce two scaled variables that we vary in our simulations: dimensionless temperature $\tilde{T}=T/T_{D}$ and effective electron–lattice coupling $\tilde{G}=2k_{F}/q_{D}$. The coupling is adjusted by varying the Fermi wavevector $k_{F}$ (energy) of the electron while maintaining the underlying lattice structure constant (with a fixed Debye wavevector $q_{D}$). This ensures that our variables $\tilde{T}$ and $\tilde{G}$ are independent, a premise supported by the evidence from our simulations.
@@ III.1 Phase Diagram
Our central finding is presented in Figure 3 showing the dynamical data classified using the ML-based clustering algorithm explained above and as detailed in Appendix A as the temperature $\tilde{T}$ and coupling strength $\tilde{G}$ are varied. Three distinct phases are identified as labeled by the differently colored regions. We want to emphasize that the term “phase” is used here to refer to regimes of different dynamical behavior, not in the thermodynamical sense. There are no sharp boundaries between these phases; the changes are gradual rather than true phase transitions. This fact is highlighted in Figure 3 by the different sizes of the points, representing the level of agreement within the ensemble of studied wavepackets for the given parameters.
@@
[…]
Figure 3: Phase diagram of LSCO in the dynamic potential field. The phase diagram was derived using a machine learning-based clustering algorithm to analyze time series data of the wavefunction evolution within systematically varied deformation potentials. This analysis involved variations in temperature $\tilde{T}$ and effective coupling $\tilde{G}$; every point corresponds to one unique configuration. The following three clusters were identified: (I) refractive scattering region (green), (II) diffraction behavior (blue), and (III) short-time localization at high temperatures (red). The size of the points indicates the level of agreement across an ensemble of different wavefunction data sets.
[…]
The final phase (Phase III) identified by the ML-clustering is associated with highly nonperturbative ($\tilde{K}\lesssim 1$) electron–lattice interaction, primarily existing in the parameter space where electron dynamics can be considered as coherent ($\tilde{\lambda}\lesssim 1$). Therefore, wave interference and diffraction effects are important because the electron wavelength is larger than the shortest length scale of the deformation potential. Notably, this phase begins at low temperature as $\tilde{T}\approx 0.5$ while extending to very high temperatures across the range of $\tilde{T}$ investigated.
In Phase III, as depicted in the right column of Figure 4, an initial wavepacket encounters significant scattering from a strong deformation potential, initially causing diffusive behavior akin to that seen in Phase II. However, wavepacket spreading eventually ceases due to quantum interference effects, signifying an onset of localization. Nevertheless, the random fluctuations introduced by the motion of the lattice slowly but surely scramble the quantum interference required for the long-term confinement of the wavepacket, resulting in the transient nature of this localization (for capturing the essential aspects of this phenomenon, see, e.g., Ref. [9]).
@@
To achieve a more comprehensive understanding, we adopt a static potential approximation, wherein the temporal aspect of the lattice deformation field of Equation (2) is neglected, effectively frozen into its original configuration. Within this frozen potential framework, we carry out an analysis analogous to that of the evolving deformation potential as above. This is outlined in Appendix B. Employing the same cluster classification, we categorize the data of an electron wavepacket evolution under a static deformation potential, yielding a phase diagram similar to its dynamic counterpart in Figure 3, albeit with slightly sharper phase boundaries. This comparison validates treating the deformation potential as predominantly static, particularly in Phases I and II. On the contrary, in Phase III, the static deformation potential results in full Anderson localization of the wavepacket that proves transient when the deformation potential undergoes morphing and undulation over time, as further elucidated in the subsequent analysis.
@@ III.2 Transient Localization
In this section, we delve deeper into the nature of transient localization induced by lattice vibrations taking place within Phase III. At timescales shorter than the characteristic timescale of $2\pi/\omega_{D}\sim 100\,\textrm{fs}$, lattice vibrations mimic a static, internal disorder field, triggering the onset of Anderson localization. Therefore, in reference to the dynamical field where the motion of the lattice disrupts the process of Anderson localization, we are also exploring the localization behavior of a wavepacket within a frozen potential approximation. In both cases, we quantify the level of localization by studying the time-dependent IPR, denoted as $\alpha(t)$ in Equation (7), and introducing a subsidiary product of the MSD, called the instantaneous diffusivity
@@
where $\beta(t)$ is defined in Equation (6).
@@
consistent with the definition used in Ref. [24]. Figure 5 shows the inverse scattering rate $1/\tau$ for both the cases of frozen (left panel (right panel) deformation potential. Overall, the analysis of the scattering rate supports the ML classification underlying the phase diagram shown in Figure 4. Both scenarios exhibit a notably high inverse scattering rate within Phase III of the phase diagram (upper right corner), indicating significant constraints on carrier mobility, as expected in the context of Anderson localization. This effect is more prominent in the frozen potential approximation than in the case of the morphing deformation potential, underlining the fact that the dynamics of the deformation potential continuously disrupt short-lived localization attempts. Moreover, we observe that the contour lines of the inverse scattering time in Figure 5 closely resemble the phase boundaries seen in Figure 4.
Figure 5: Inverse scattering rate $1/\tau$ for electrons in LSCO in a (a) frozen and (b) dynamic deformation potential on a normalized scale. Both scenarios exhibit a high inverse scattering rate region in the upper left corner (Phase III), suggesting strong constraints on carrier mobility. This region is markedly more prominent in the frozen potential scenario, as the lattice vibrations in the dynamic potential continuously disrupt short-lived attempts at localization. The shape of counter lines closely resembles the dynamical phase transition lines depicted in Figure 4.
@@
A deeper insight into the emergence of transient localization is obtained by investigating the IPR of the wavepacket evolving over time. In the left panel of Figure 6, we show the evolution of the IPR in Phase III for both static (blue) and dynamic (green) deformation potentials. In the approximation where the deformation potential remains static at its initial state, the IPR stabilizes at a certain value ($\beta\sim 0.3$) after an initial decrease, heralding the Anderson localization of the wavepacket. Similarly, when the wavepacket is subject to a dynamic deformation potential, a form of Anderson localization occurs. This localized state is eventually disrupted by lattice motion, leading to a temporary delocalization followed by a brief relocalization before being disintegrated again by potential evolution. This cyclical process of plateauing IPR seen in the left panel of Figure 6 epitomizes the birth and demise of the Anderson localization due to lattice vibrations.
@@
We would like to point out that the dynamics of short wavelength components of the deformation potential can significantly influence localization within a time window shorter than the characteristic time $\omega_{D}/2\pi$. For example, the left panel of Figure 6 demonstrates an instance of dynamically enhanced localization: the random initial configuration of the potential creates valleys (mountains) that quickly move towards (away) the wavefunction, causing boosted localization. Vice versa, these types of small potential variations in time can also lead to weaker localization compared to the static potential case in the first localization plateau seen in the left panel of Figure 6.
@@
To provide a full picture, we also present the evolution of the IPR measure for Phase I and II in the right and left panels of Figure 6, respectively. Neither phase exhibits any signs of localization, contrasting with the behavior observed in Phase III. Moreover, the overlapping curves of the static and dynamic potentials further support the earlier assertion that the deformation potential can be effectively approximated as a static entity in Phase I and II. In Phase II, the IPR shows rapid exponential decay, quickly approaching the ergodic (fully delocalized) limit of $\beta\sim 0$. This behavior resembles that of a system with a high density of impurities characterized by Gaussian statistics. Similarly, in Phase I, the IPR decreases towards the ergodic limit, albeit at a slower rate, displaying subtle oscillation. The slower, non-exponential decay can be attributed to the weak, refractive nature of wavepacket scattering, in conjunction with quantum coherence and interference effects.
Figure 6: Normalized inverse participation ratio $\beta(t)$ of the wavepacket as a function of time across three distinct dynamical regions identified in the machine learning-based phase diagram shown in Figure 4. Figure contrasts the behaviors under static (blue line) and dynamic (green line) deformation potentials. In Phase III (left), brief localization periods (indicated by dashed lines) are broken by lattice motion (indicated by arrows), while full localization, indicated by the saturation of $\beta(t)$, occurs within the frozen potential approximation. In Phase II (middle) and Phase I (right), the decay of $\beta(t)$ follows a relatively exponential trend without signs of localization as observed in Phase III. Notably, the decay is faster in Phase I when compared to Phase II, where small oscillations linked to quantum interference are present.
@@ IV Conclusion and Future Directions
Quantum acoustics opens up an unexplored pathway to investigating the intricate Fröhlichian electron–lattice interaction inaccessible with the standard methods of perturbation theory. We take the concept further by treating it not just as a dual perspective on lattice vibrations, but as a versatile tool in ascendance: a time-dependent, nonperturbative approach for electron–lattice interaction in coordinate space. Moreover, in the quasi-classical limit of the coherent state formalism, the quantum-acoustical way unveils the dynamics of electrons navigating through an internal lattice disorder field undulating and propagating in time.
@@
Our study, supported by machine learning, explores the parameter space characterized by temperature and effective coupling, focusing on the paradigmatic strange metal LSCO, known for its two-dimensional transport behavior. However, the presented method can be readily extended to variations in any set of material parameters—potentially augmented by density functional simulations as they are not necessarily independent—and generalized to electron–lattice dynamics in three dimensions. Therefore, our work not only lays the groundwork for uncovering hidden realms in electron–lattice interaction but is also a testament to designing materials with customized features by employing machine learning techniques from a dynamics perspective.
@@ Appendix A Clustering Analysis
    Update Step: Update the centroids using $\mathcal{S}$ according to Equation (10).
It is easy to see that these steps cannot increase the in-set variance; the algorithm is guaranteed to converge. However, there is an important caveat in that the objective function $\mathcal{L}(\mathcal{S})$ defined in Equation (9) is non-convex and the algorithm can therefore converge to a local optimum that is not the global minimum. This problem can be mitigated by an intelligent choice of initialization scheme, most commonly $k$-means++ [62], and by running the algorithm multiple times and taking the clustering with the minimal $\mathcal{L}(\mathcal{S})$ post-convergence [27]. An example run of $k$-means on toy data in $\mathbb{R}^{2}$ is shown in Figure A.1.
Figure A.1: Toy example to illustrate a run of the $k$-means algorithm with hyperparameter $k=3$ on example data in $\mathbb{R}^{2}$ (black dots). (a) Initial centroids are chosen by an initialization scheme (here: random initialization). (b) Assignment step of the first iteration. (c) Update step of the first iteration. (d) Final result. The algorithm converges after two consequent iterations yield sufficiently similar clusters (here the $10$th iteration).
@@
When working with time series data $x(t)=(x^{t_{1}},x^{t_{2}},\ldots,x^{t_{M}})$, $x^{t_{i}}\in\mathbb{R}^{d}$, one could naively embed $x(t)$ in the space $\mathbb{R}^{M\times d}$ (here, $M=50$ and $d=2$) equipped with the Euclidean metric and run $k$-means exactly as described above. However, this approach has significant limitations. It only works for sequences of equal length, and more critically, the Euclidean metric is unable to account for the temporal nature of the sequences. By operating solely on elements with the same time indices, it ignores potential temporal misalignment between sequences. Consequently, even time series with similar features can result in a large metric distance if their phases differ. This makes it an ineffective measure for determining similarity. Dynamic time warping (DTW), introduced by Sakoe and Chiba [51], has been employed to overcome these challenges by identifying the most temporally appropriate pairs of time indices for comparison [63]. Formally,
@@
where the time index pairs $\pi=\{\pi_{0},\pi_{1},\ldots,\pi_{K}\}$ satisfy certain (boundary) conditions that we will not elaborate on for brevity (for an extensive introduction to the topic, see, e.g., Ref. [64]). One thus ends up with a new, warped time path for the comparison of the time series, as is illustrated in Figure A.2.
Figure A.2: Illustration of the difference between the Euclidean distance and Dynamic Time Warping (DTW) distance for two example MSD sequences from our data set. While the Euclidean distance compares values at corresponding time steps, DTW temporally aligns the sequences, providing a more robust distance measure for time series data.
DTW is not a metric, because $\text{DTW}(x(t),y(t))=0$ does not imply $x(t)=y(t)$ and it does not obey the triangle inequality. Nevertheless, the Fréchet mean, which generalizes Equation (10) to any metric space, is used to compute a centroid-like representation, commonly referred to as barycenter:
@@
$\sigma_{t}=\sqrt{\frac{1}{T}\int_{0}^{T}\left(x(t)-\mu_{t}\right)^{2}\,\text{d}t}$
@@
The parameter $k$, i.e., the number of clusters, has to be set before performing the optimization Equation (9). To find a suitable $k$ for our analysis, we run the $k$-means algorithm multiple times with varying values of $k$ ranging from 1 to 10. Figure A.3 shows a plot of the sum of squared distances of samples to their nearest cluster center (inertia) against different values of $k$. The point where the curve bends or forms an elbow is considered an indicator of the natural number of clusters in the underlying data [27]; additional clusters beyond this point do not substantially improve the model’s representation of the different dynamical regimes of the system.
Choosing $k$
Figure A.3: Elbow plot depicting the sum of squared distances of samples to their nearest cluster center for varying values of $k$ for the (a) frozen and (b) dynamic lattice vibration field scenario. The elbow point (indicated by a red star), where the curve shows a distinct bend, suggests the optimal number of clusters for our analysis, as increasing the number of clusters beyond this value yields diminishing returns in terms of model improvement.
@@
To mitigate the effect of statistical fluctuations of the lattice vibration field on the clustering of the electron dynamics, we perform the optimization independently on an ensemble of 10 wavefunction data sets, each generated using deformation potentials with different random initializations. The different clustering results are then consolidated into one phase diagram by taking the consensus of the ensemble, with the size of the dots in Figure 3 indicating the degree of consensus.